A researcher examined whether the time of day someone exercises affects memory retention in college courses.  Participants were assigned to one of three exercise groups: morning, afternoon, evening, and their performance on a memorization task was measured.  This data is below:

Morning Afternoon Evening
6 4 7
7 5 6
8 6 8
5 4 6
6 5 5
Mean = 6.40 Mean = 4.80 Mean = 6.40
s = 1.14 s =.84 s = 1.14

Assuming the researcher wants to know whether the performance was different in these groups (alpha = .05).  What is the Total Sum of Squares (SST)?  Round to two decimal places.
Answer #1

First be need to find grand mean so

\bar{x}=\frac{5\cdot \left (6.40+4.80+6.40 \right )}{15}=5.87

Now,

SS_{between}=n_{1}(\bar{x}_{1}-\bar{x})^{2}+n_{2}(\bar{x}_{2}-\bar{x})^{2}+n_{3}(\bar{x}_{3}-\bar{x})^{2}=5\cdot (6.40-5.87)^{2}+5\cdot (4.80-5.87)^{2}+5\cdot (6.40-5.87)^{2}=8.53

and

SS_{within}=(n_{1}-1)s^{2}_{1}+(n_{2}-1)s^{2}_{2}+(n_{3}-1)s^{2}_{3}=13.22

Therefore

SS_{total}=SS_{between}+SS_{within}=21.75